# A quasilinear Keller-Segel model with saturated discontinuous advection

### v2

@@ 4.1. Entropy method
$\leq$ $\;-\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}+\sigma}{2}\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{|\nabla\rho_{\varepsilon}|}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{|\nabla\rho_{\delta}|}\right)_{+}\right]^{2}\,{\rm d}x\,{\rm d}t$
$-\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}+\sigma}{2}\left|\nabla\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right|^{2}\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t$
@@
$\;\;\;-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{|\nabla\rho_{\varepsilon}|}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{|\nabla\rho_{\delta}|}\right)_{+}\right]^{2}\,{\rm d}x\,{\rm d}t$
$\;\;\;-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\nabla\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right|^{2}\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t$
$\;\;\;+R_{1}+R_{2}+{I_{1}+I_{2}}+E_{\varepsilon,\delta},$ (38)
where, $R_{1}$, $R_{2}$, $I_{1}$, $I_{2}$ and $E_{\varepsilon,\delta}$ vanish when $\sigma,\varepsilon,\delta\rightarrow 0$. Now, we need to prove
${\int_{\mathbb{R}^{d}}(\rho_{\delta}(T)-\rho_{in,\delta})\phi_{R}\,{\rm d}x}\rightarrow 0$ when $\sigma,\varepsilon,\delta\rightarrow 0$. For this, we test (3.2) with $\phi_{R}$ to get
@@
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\rho_{\delta}\nabla\phi_{R}\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\delta}\phi_{R}\,{\rm d}x\,{\rm d}t.$ (39)
Since $|\nabla\phi_{R}|\leq\frac{C}{R(\sigma)}\rightarrow 0$ and $\delta\rightarrow 0$ when $\sigma\rightarrow 0$, we have that
@@
$\lim_{\sigma\rightarrow 0}\left\lvert\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\nabla\phi_{R}\,{\rm d}x\,{\rm d}t\right\rvert\leq\lim_{\sigma\rightarrow 0}\frac{C}{R(\sigma)}\|\nabla\sqrt{\rho_{\delta}}\|_{2}\|\rho_{\delta}\|_{1}=0.$
The viscosity terms in (4.1) also vanish when $\sigma\rightarrow 0$, and therefore we get that
$\lim_{\sigma\rightarrow 0}\left(\int_{\mathbb{R}^{d}}\rho_{\delta}\varphi\,{\rm d}x\right)(T)=\lim_{\sigma\rightarrow 0}\int_{\mathbb{R}^{d}}\rho_{{\rm in},\delta}\varphi\,{\rm d}x.$
Finally, taking the limit $\sigma,\varepsilon,\delta\rightarrow 0$ in (4.1) yields
@@
$\lim_{\epsilon,\delta\rightarrow 0}\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{|\nabla\rho_{\varepsilon}|}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{|\nabla\rho_{\delta}|}\right)_{+}\right]^{2}\phi_{R}\,{\rm d}x\,{\rm d}t=0,$ (40a)
@@
$\lim_{\epsilon,\delta\rightarrow 0}\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\nabla\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right|^{2}\cdot$
@@
$\;\;\;\;\;\;\;\;\;\;\;\;\cdot\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\phi_{R}\,{\rm d}x\,{\rm d}t=0.$ (40b)
4.2. Identification of the limit of $\nabla\rho_{\varepsilon}$
Now, from (40a) is easy to see that
$\left(1-\frac{\chi(\rho_{\varepsilon}+\varepsilon)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\rightarrow\overline{G}\quad\mbox{a.e. }x\in[0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\}.$
And on the other hand, from (40b)
$\frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}+\varepsilon}\rightarrow w\quad\mbox{a.e. in }[0,T]\times\mathbb{R}^{d}\setminus\{\{\rho=0\}\cup\{\overline{G}=0\}\}.$
Notice, moreover, that if $\left(1-\frac{\chi(\rho_{\varepsilon}+\varepsilon)}{|\nabla\rho_{\varepsilon}|}\right)_{+}$ converges almost everywhere in $[0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\}$ then
$\frac{\left|\nabla\rho_{\varepsilon}\right|}{\rho_{\varepsilon}+\varepsilon}<\infty\quad\mbox{a.e. in }[0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\},$
and we have
@@ 4.2. Identification of the limit of $\nabla\rho_{\varepsilon}$
$(\rho_{\varepsilon}+\varepsilon)^{3}\frac{\left|\nabla\rho_{\varepsilon}\right|}{\rho_{\varepsilon}+\varepsilon}\left(\frac{\left|\nabla\rho_{\varepsilon}\right|}{\rho_{\varepsilon}+\varepsilon}+\chi\right)\left(1-\frac{\chi(\rho_{\varepsilon}+\varepsilon)}{|\nabla\rho_{\varepsilon}|}\right)\frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}+\varepsilon}\rightarrow\overline{w},$
almost everywhere in $[0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\}$, or, which is equivalent,
@@
$\left(|\nabla\rho_{\varepsilon}|^{2}-(\rho_{\varepsilon}+\varepsilon)^{2}\chi^{2}\right)_{+}\nabla\rho_{\varepsilon}\rightarrow\overline{w},$

### v3

@@ 4.1. Contractivity of the relative entropy
$\leq$ $\;-\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}+\sigma}{2}\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{|\nabla\rho_{\varepsilon}|}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{|\nabla\rho_{\delta}|}\right)_{+}\right]^{2}\,{\rm d}x\,{\rm d}t$
$-\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}+\sigma}{2}\left|\nabla\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right|^{2}\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t$
@@
$\;\;\;-\frac{1}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{|\nabla\rho_{\varepsilon}|}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{|\nabla\rho_{\delta}|}\right)_{+}\right]^{2}\,{\rm d}x\,{\rm d}t$
$\;\;\;-\frac{1}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\nabla\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right|^{2}\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\hskip-7.11317pt+\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t$
$\;\;\;+R_{1}+R_{2}+E_{\varepsilon,\delta,\sigma},$ (38)
where $R_{1}+R_{2}\to 0$ as $\sigma\to 0$, for every $\varepsilon,\delta>0$, while $E_{\varepsilon,\delta,\sigma}\leq C(\varepsilon+\delta)$ for every $\sigma>0$.
Now, we are going to find a bound for $\int_{\mathbb{R}^{d}}(\rho_{\delta}(T)-\rho_{in,\delta})\,{\rm d}x$}. This is done in a rigorous way by employing a cutoff $\phi_{R}=\phi_{R}(x)$ as test function in (19). The cutoff $\phi_{R}$ has the following properties:
[…]
We obtain
@@
$-\delta\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\rho_{\delta}\nabla\phi_{R}\,{\rm d}x\,{\rm d}t-\delta\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\delta}\phi_{R}\,{\rm d}x\,{\rm d}t.$ (39)
Since $|\nabla\phi_{R}|\leq\frac{C}{R}$ we have that
@@
$\left\lvert\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\nabla\phi_{R}\,{\rm d}x\,{\rm d}t\right\rvert\leq\frac{C}{R}\|\nabla\sqrt{\rho_{\delta}}\|_{2}\|\rho_{\delta}\|_{1}^{1/2}{\leq\frac{C}{R}}.$
A similar estimate can be obtained for the viscosity terms in (4.1), and therefore we get that
${\left|\left(\int_{\mathbb{R}^{d}}\rho_{\delta}{\phi_{R}}\,{\rm d}x\right)(T)-\int_{\mathbb{R}^{d}}\rho_{{\rm in},\delta}{\phi_{R}}\,{\rm d}x\right|\leq\frac{C}{R}+C\delta}.$
Taking the limit $R\to\infty$ in the above estimate leads to
${\left|\left(\int_{\mathbb{R}^{d}}\rho_{\delta}\,{\rm d}x\right)(T)-\int_{\mathbb{R}^{d}}\rho_{{\rm in},\delta}\,{\rm d}x\right|\leq C\delta}.$
Hence, taking the (superior) limit $\sigma\rightarrow 0$ in (4.1) and applying Fatou’s Lemma yield
@@
$\limsup_{\sigma\to 0}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{|\nabla\rho_{\varepsilon}|}\right)_{+}\hskip-11.38109pt-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{|\nabla\rho_{\delta}|}\right)_{+}\right]^{2}\,{\rm d}x\,{\rm d}t\leq H_{rel}[\rho_{in,\varepsilon}|\rho_{in,\delta}]+C(\varepsilon+\delta),$
@@
$\limsup_{\sigma\rightarrow 0}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\nabla\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right|^{2}\cdot$
@@
$\;\;\;\;\;\;\;\;\;\;\;\;\cdot\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t\leq H_{rel}[\rho_{in,\varepsilon}|\rho_{in,\delta}]+C(\varepsilon+\delta).$
We need now to compute the a.e. limit as $\sigma\to 0$ of the integrands in the above relations. It holds:
@@
$\mbox{On }\{\rho_{\varepsilon}>0\}:\leavevmode\nobreak\ \leavevmode\nobreak\ \left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{|\nabla\rho_{\varepsilon}|}\right)_{+}\to\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{ as }\sigma\to 0,$
$\mbox{on }\{\rho_{\varepsilon}=0\}:\leavevmode\nobreak\ \leavevmode\nobreak\ \left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{|\nabla\rho_{\varepsilon}|}\right)_{+}=\left(1-\frac{\chi\sigma}{|\nabla\rho_{\varepsilon}|}\right)_{+}=0,$
since $\nabla\rho_{\varepsilon}=0$ a.e. on $\{\rho_{\varepsilon}=0\}$. This means that
$\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{|\nabla\rho_{\varepsilon}|}\right)_{+}\to\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}{\bf 1}_{\{\rho_{\varepsilon}>0\}}\quad\mbox{ a.e.\leavevmode\nobreak\ in $\mathbb{R}^{d}\times(0,T)$ as }\sigma\to 0.$
Let us now consider
$\nabla\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)=\frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}+\sigma}-\frac{\nabla\rho_{\delta}}{\rho_{\delta}+\sigma}.$
It holds
$\mbox{On }\{\rho_{\varepsilon}>0\}:\leavevmode\nobreak\ \leavevmode\nobreak\ \frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}+\sigma}\to\frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}}\quad\mbox{ as }\sigma\to 0,$
$\mbox{on }\{\rho_{\varepsilon}=0\}:\leavevmode\nobreak\ \leavevmode\nobreak\ \frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}+\sigma}=\frac{\nabla\rho_{\varepsilon}}{\sigma}=0,$
again because $\nabla\rho_{\varepsilon}=0$ a.e. on $\{\rho_{\varepsilon}=0\}$. It follows
$\nabla\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\to\frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}}{\bf 1}_{\{\rho_{\varepsilon}>0\}}-\frac{\nabla\rho_{\delta}}{\rho_{\delta}}{\bf 1}_{\{\rho_{\delta}>0\}}\quad\mbox{ as }\sigma\to 0.$
Hence we get
@@
$\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}{\bf 1}_{\{\rho_{\varepsilon}>0\}}-\left(1-\frac{\rho_{\delta}}{|\nabla\rho_{\delta}|}\right)_{+}{\bf 1}_{\{\rho_{\delta}>0\}}\right]^{2}\,{\rm d}x\,{\rm d}t\leq H_{rel}[\rho_{in,\varepsilon}|\rho_{in,\delta}]+C(\varepsilon+\delta),$
$\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left|\frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}}{\bf 1}_{\{\rho_{\varepsilon}>0\}}-\frac{\nabla\rho_{\delta}}{\rho_{\delta}}{\bf 1}_{\{\rho_{\delta}>0\}}\right|^{2}\cdot$
$\;\;\;\;\;\;\;\;\;\;\;\;\cdot\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}{\bf 1}_{\{\rho_{\varepsilon}>0\}}+\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}{\bf 1}_{\{\rho_{\delta}>0\}}\right]\,{\rm d}x\,{\rm d}t\leq H_{rel}[\rho_{in,\varepsilon}|\rho_{in,\delta}]+C(\varepsilon+\delta).$
Therefore by taking the limit $\varepsilon,\delta\to 0$ in the above estimates we obtain
$\lim_{\epsilon,\delta\rightarrow 0}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}{\bf 1}_{\{\rho_{\varepsilon}>0\}}-\left(1-\frac{\rho_{\delta}}{|\nabla\rho_{\delta}|}\right)_{+}{\bf 1}_{\{\rho_{\delta}>0\}}\right]^{2}\,{\rm d}x\,{\rm d}t=0,$ (40a)
$\lim_{\epsilon,\delta\rightarrow 0}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left|\frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}}{\bf 1}_{\{\rho_{\varepsilon}>0\}}-\frac{\nabla\rho_{\delta}}{\rho_{\delta}}{\bf 1}_{\{\rho_{\delta}>0\}}\right|^{2}\cdot$
$\;\;\;\;\;\;\;\;\;\;\;\;\cdot\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}{\bf 1}_{\{\rho_{\varepsilon}>0\}}+\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}{\bf 1}_{\{\rho_{\delta}>0\}}\right]\,{\rm d}x\,{\rm d}t=0.$ (40b)
4.2. Identification of the limit of the flux
Our goal is to show that
$\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\rightarrow\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\quad\mbox{a.e. in }(0,T)\times\mathbb{R}^{d}.$
The above a.e. convergence relation will then allow us to identify the weak $L^{2}$ limit of the term $\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}$. By Lemma 4, we know that
Given mass conservation it follows that $\rho_{\varepsilon}\to\rho$ strongly in $L^{p}((0,T)\times\mathbb{R}^{d})$ for every $p>1$. Putting this fact together with the uniform moment bounds for $\rho_{\varepsilon}$ we deduce that
$\rho_{\varepsilon}\rightarrow\rho\quad\mbox{strongly in }L^{1}((0,T)\times\mathbb{R}^{d}).$
From the above convergence property and (40a) one easily obtains
$\lim_{\epsilon,\delta\rightarrow 0}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho\left[\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}{\bf 1}_{\{\rho_{\varepsilon}>0\}}-\left(1-\frac{\rho_{\delta}}{|\nabla\rho_{\delta}|}\right)_{+}{\bf 1}_{\{\rho_{\delta}>0\}}\right]^{2}\,{\rm d}x\,{\rm d}t=0$
which means that the sequence $\left(\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}{\bf 1}_{\{\rho_{\varepsilon}>0\}}\right)_{\varepsilon>0}$ is Cauchy in $L^{2}(\mathbb{R}^{d}\times(0,T),\rho\,{\rm d}x\,{\rm d}t)$. Therefore, a (unique) function $\Xi\in L^{2}(\mathbb{R}^{d}\times(0,T))$ exists such that
$\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}{\bf 1}_{\{\rho_{\varepsilon}>0\}}\rightarrow\Xi\quad\mbox{strongly in }L^{2}(\mathbb{R}^{d}\times(0,T),\rho\,{\rm d}x\,{\rm d}t).$ (41)
In particular, up to subsequences it holds
$\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}{\bf 1}_{\{\rho_{\varepsilon}>0\}}\rightarrow\Xi\quad\mbox{a.e.\leavevmode\nobreak\ in }[0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\}.$ (42)
And on the other hand, from (40b) we deduce the a.e. convergence (up to subsequences) of the integrand:
$\lim_{\epsilon,\delta\rightarrow 0}\rho_{\varepsilon}\left|\frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}}{\bf 1}_{\{\rho_{\varepsilon}>0\}}-\frac{\nabla\rho_{\delta}}{\rho_{\delta}}{\bf 1}_{\{\rho_{\delta}>0\}}\right|^{2}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}{\bf 1}_{\{\rho_{\varepsilon}>0\}}+\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}{\bf 1}_{\{\rho_{\delta}>0\}}\right]=0,$
$\mbox{a.e.\leavevmode\nobreak\ in }\mathbb{R}^{d}\times(0,T)$
therefore
$\lim_{\epsilon,\delta\rightarrow 0}\left|\frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}}{\bf 1}_{\{\rho_{\varepsilon}>0\}}-\frac{\nabla\rho_{\delta}}{\rho_{\delta}}{\bf 1}_{\{\rho_{\delta}>0\}}\right|=0\quad\mbox{a.e.\leavevmode\nobreak\ on }[0,T]\times\mathbb{R}^{d}\setminus\{\{\rho=0\}\cup\{\Xi=0\}\}.$
This means that a.e. in $\mathbb{R}^{d}\times(0,T)$ the sequence $(\frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}}{\bf 1}_{\{\rho_{\varepsilon}>0\}})_{\varepsilon>0}$ is Cauchy and therefore convergent towards some function $w$: